OpenStudy (anonymous):

A complex number will, in general, have _____ fifth complex roots. Help fill in the blank? :o

OpenStudy (psymon):

Whatever the number complex root, that's how many there are. 5 5th complex, 4 4th complex, 3 3rd complex, etc.

OpenStudy (anonymous):

Oh okay that's easy! :)

OpenStudy (anonymous):

how do you find the fourth root of \[-1+i \sqrt{3}\] ?

OpenStudy (psymon):

I'm not positive, but I think we need it in polar form first. At least I know how to do it from polar form, lol.

OpenStudy (psymon):

You remember how to put it in polar form? xD

OpenStudy (anonymous):

NO D: >.< :(

OpenStudy (psymon):

Forgot since last night xD Bleh. Well, we need r(costheta + isintheta). r comes from the same old pythagorean theorem thing. \[r = \sqrt{a ^{2}+b ^{2}}\]. A is your negative 1 and b is your sqrt(3) So we can start off by getting r :3

OpenStudy (anonymous):

So it'd be 1+ 3 = 4 ?

OpenStudy (anonymous):

Square root of 3 Squared is 3 ?

OpenStudy (psymon):

Right. So you would just have sqrt of 4 which is 2 xD

OpenStudy (psymon):

So r = 2 :P Now to get the angle, we need to say that \[\tan \theta = \frac{ b }{ a }\]

OpenStudy (anonymous):

so, \[4\sqrt{2}\].. b is \[i \sqrt{3} ?\]

OpenStudy (psymon):

Whoa, wait xD Not sure where 4rt2 came from o.o a = -1 b = \[\sqrt{3}\] no i included \[r = \sqrt{(-1)^{2}+\sqrt{3}^{2}} = \sqrt{4} = 2\] This is where we are so far xD Now since a = -1 and b = sqrt 3, we need to say: \[\tan \theta =\frac{ \sqrt{3} }{ -1 }\]

OpenStudy (anonymous):

ohh, okkay !, which is \[-\sqrt{3} ??\] D:

OpenStudy (psymon):

Lol, yes xD No need to freak out, that one is correct :P So now, tangent is negative in the 2nd and 4th quadrants. So if our original rectangular says \[-1 + i \sqrt{3}\], then which quadrant is that,2 or 4?

OpenStudy (anonymous):

2 ?

OpenStudy (psymon):

Yep :p The rectangular form basically says left one and up rt(3) xD So now, in the 2nd quadrant, any idea where tan(theta) = -sqrt(3)?

OpenStudy (anonymous):

1 quadrant ?

OpenStudy (anonymous):

._o

OpenStudy (psymon):

No, no, no, it's in the 2nd quadrant xD That's what I was just basically asking :P  Because tangent is the division of y/x, the only way it can be negative is if it's in a quadrant where either x or y, but not both, are negative. So on that circel, you can see that quadrant 2 and quadrant 4 are the places where you ciuld have a negative tangent.

OpenStudy (psymon):

If you're still alive xD